OpenStudy (anonymous):

PLEASE HELP Given integral -1 to 1 f(x) dx = 0 and integral 0 to 1 f(x)dx = 5 find

OpenStudy (gorv):

what we have to find??

OpenStudy (anonymous):

integral from -1 to 0 f(x) dx

OpenStudy (anonymous):

integral from 0 to 1 f(x)dx -integral from -1 to 0 f(x)dx integral from -1 to 1 3f(x)dx integral form 0 to 1 3f(x)dx

OpenStudy (gorv):

ok u have \[\int\limits_{-1}^{1}\]

OpenStudy (gorv):

\[\int\limits_{-1}^{1}=\int\limits_{-1}^{0}+\int\limits_{0}^{1}\]

OpenStudy (gorv):

\[\int\limits_{-1}^{1}-\int\limits_{0}^{1}=\int\limits_{-1}^{0}\]

OpenStudy (gorv):

can u do it now @halrenw